# Whose Emotions and Moral Sentiments Do Language Models Reflect?

### v1

@@ Abstract
Language models (LMs) are known to represent the perspectives of some social groups better than others, which may impact their performance, especially on subjective tasks such as content moderation and hate speech detection. To explore how LMs represent different perspectives, existing research focused on positional alignment, i.e., how closely the models mimic the opinions and stances of different groups, e.g., liberals or conservatives. However, human communication also encompasses emotional and moral dimensions. We define the problem of affective alignment, which measures how LMs’ emotional and moral tone represents those of different groups. By comparing the affect of responses generated by 36 LMs to the affect of Twitter messages, we observe significant misalignment of LMs with both ideological groups. This misalignment is larger than the partisan divide in the U.S. Even after steering the LMs towards specific ideological perspectives, the misalignment and liberal tendencies of the model persist, suggesting a systemic bias within LMs. [^1] Code and data are available at https://github.com/zihaohe123/llm-affective-alignment.
@@ 1 Introduction
The capacity of language models (LMs) to generate human-like responses to natural language prompts has led to new technologies that support people on cognitive tasks requiring complex judgements. However, researchers found that LMs inherit biases [^2] Throughout this paper, we use “bias” to refer to a systematic statistical tendency, rather than unfairness or prejudice. from humans, as their views are shaped by online users who produced the pretraining data, feedback from crowdworkers during Reinforcement Learning from Human Feedback (RLHF) process Ouyang et al. (2022), and potentially, the decisions made by the model developers themselves Santurkar et al. (2023). In subjective tasks, such as hate speech detection Hartvigsen et al. (2022), content moderation He et al. (2023), and legal judgement Jiang and Yang (2023), these biases may show up as LMs adopting the perspectives of one group while excluding others. This may lead to undesirable consequences in downstream applications, ranging from negative individual user experiences with LM-driven interfaces to societal level division and polarization.
@@
To examine how LMs represent differences in perspectives of different groups, existing research has looked at positional alignment: how closely the opinions, stances, or positions exhibited by a model mirror those of different social groups Santurkar et al. (2023); Durmus et al. (2023). Using multi-choice survey questions, researchers have demonstrated that language models are misaligned with the US population and represent the perspectives of some demographic groups better than others.
@@
Figure 1: The framework for evaluating affective alignment of LMs. We first prompt LMs to generate tweets on a topic using default prompting or steered prompting. The distributions of emotions and moral sentiments of LM-generated tweets are then compared to that of human-written tweets. Affective alignment is measured as one minus the Jensen-Shannon distance (JSD) between the two distributions.
@@
However, positional agreement captures just one aspect of alignment. Human communication also carries cues to emotions and moral sentiments–collectively referred to as affect–which are integral to social interaction and cohesion Graham et al. (2009); Iyengar et al. (2019); Mokhberian et al. (2020). How LMs represent affect plays an important role in their performance in downstream tasks, especially in subjective tasks. Consider how an LM facilitating online discussions may handle the following comment: “Wearing a mask is a personal choice, not a public responsibility.” An LM aligned with conservatives would not flag this comment as it prioritized the moral sentiments of liberty and authority typically associated with conservatives Doğruyol et al. (2019). However, this comment may evoke negative reactions from liberals, as it goes against their deeply-held values of care and fairness, and thus an LM aligned with liberals are likely to flag it. This motivates us to ask:
@@
Whose affect, i.e., moral and emotional tone, do language models reflect?
@@
Our contributions. We define the problem of affective alignment, which measures how closely the emotional or moral tone of the model matches what people express in similar circumstances. To represent human affect, we study two datasets of Twitter messages about contentious issues such as the COVID-19 pandemic and abortion. To analyze differences between groups, we disaggregate users based on their detected political ideology as liberal or conservative. [^3] In this work we focus on the liberals and conservatives within the context of U.S. politics, but our framework should naturally generalize to other demographic groups. We prompt 36 LMs of varying size, from millions to billions of parameters, to generate statements on contentious topics, such as “COVID-19 mask mandates” and “abortion rights and access” and then compare the affect (emotions and moral sentiments) in model-generated responses to that expressed by Twitter users belonging to different ideological groups.
@@
We first assess the models’ default affective alignment, based on the responses they generate to default prompts that do not include information about a target demographic (persona). Our findings suggest that LMs show significant misalignment in affect with either ideological group, and the differences are larger than the ideological divide between partisan groups on Twitter. Moreover, consistent with prior findings Santurkar et al. (2023); Perez et al. (2022); Hartmann et al. (2023), all LMs exhibit liberal tendencies on topics related to COVID-19.
@@
Next, we assess LMs’ affective alignment after steering, i.e., when we provide additional context in the prompt to generate texts from the perspective of liberals or conservatives. The results reveal that steering can better align the affect of the models with the target group for most instruction-tuned LMs. However, even after steering, the models remain misaligned. In addition, the liberal tendencies of LMs cannot be mitigated simply by steering.
@@
We believe that a deep analysis of the affect expressed by existing LMs is crucial for building AI systems for greater social good. To the best of our knowledge, our work is the first to systematically assess the affective alignment of LMs, which highlights the unequal affective representations of different ideological groups in current LMs. We hope that our framework can help guide future research in better understanding LMs’ representativeness of people from diverse backgrounds on an emotional and moral level.
@@ 2 Measuring Affective Alignment
Figure 1 shows the overall framework. We first prompt LMs to generate tweets on some topic using default prompting or steered prompting. We use state-of-the-art classifiers to measure emotions and moral sentiment of both LM-generated tweets and human-written tweets. The affective alignment is measured as one minus the Jensen-Shannon distance (JSD) between the distributions of affect.
@@ 2.1 Online Sociopolitical Discourse
To represent human affect empirically, we compile two Twitter datasets on sociopolitical discourse—COVID-19 Tweets and Roe v. Wade Tweets—and identify fine-grained sub-topics related to both issues. We estimate the ideological leanings (liberals vs. conservatives) of users based on the political bias of the news outlets they share. See Appendix A.1 for details.
@@ 2.2 Text Generated by LMs
We categorize LMs into two groups: 1) base LMs Ouyang et al. (2022); Touvron et al. (2023); Almazrouei et al. (2023); Jiang et al. (2023); Conover et al. (2023); Chung et al. (2022); Zheng et al. (2023), that have been pretrained with the causal language modeling objective using data from the Internet; 2) instruction-tuned LMs Ouyang et al. (2022); Touvron et al. (2023); Almazrouei et al. (2023); Jiang et al. (2023); Conover et al. (2023); Chung et al. (2022); Zheng et al. (2023), that have been further finetuned on instructions following and with RLHF for better alignment with human values. To guide the LM to generate open-ended responses on different topics, we use opinion-eliciting prompts. To measure affect representativeness of different configurations of models, we consider the following two settings:
[…]
@@
The idea for these two kinds of prompting is inspired by previous works Santurkar et al. (2023); Durmus et al. (2023). To mitigate the effect of the model’s sensitivity to the specific wording in a prompt, we craft 10 different prompts for the base LMs and instruction-tuned LMs, using default prompting and steered prompting, respectively (Table 3 in Appendix). For each fine-grained topic, we generate 2,000 responses, using 2,000 prompts randomly sampled (with replacement) from the 10 candidate prompts. For more details on the generation process, please see Appendix A.2.
@@ 2.3 Measuring Affect
Detecting Emotions. Emotions are a powerful element of human communication vanKleef et al. (2016). To detect emotions, we use SpanEmo Alhuzali et al. (2021), fine-tuned on top of BERT Devlin et al. (2019) on the SemEval 2018 1e-c data Mohammad et al. (2018), which is specifically curated from Twitter and widely recognized as a benchmark for emotion detection on social media. SpanEmo learns the correlations among the emotions and achieves a micro-F1 score of 0.713 on this dataset, outperforming several other baselines and achieving the state-of-the-art in detecting emotions on Twitter data. We measure the following emotions: anticipation, joy, love, trust, optimism, anger, disgust, fear, sadness, pessimism and surprise. The model returns a score giving the confidence that a tweet expresses an emotion. We average scores over all tweets with that emotion.
Detecting Moral Language. Moral Foundations Theory Haidt et al. (2007) posits that individuals’ moral perspectives are a combination of a set of foundational values. These moral foundations are quantified along five dimensions: dislike of suffering (care/harm), dislike of cheating (fairness/cheating), group loyalty (loyalty/betrayal), respect for authority and tradition (authority/subversion), and concerns with purity and contamination (purity/degradation). These moral dimensions are crucial for understanding the values driving liberal and conservative discourse. We use DAMF Guo et al. (2023b) for morality detection, which is finetuned on top of BERT Devlin et al. (2019) on three Twitter datasets (including COVID-19 tweets and abortion-related tweets studied in this paper) and one news article dataset. The large amount and the variety of topics in the training data helps mitigate the data distribution shift during inference. The model returns a value indicating a confidence that a tweet expresses a moral foundation. We average scores over all tweets with that moral foundation.
On the accuracy of measuring affect. Please refer to Limitations.
@@ 2.4 Measuring Alignment
Let us represent an LM as $f$ and a group of humans as $g$. We aim to measure affective alignment $S^{T}(f,g)$ between the LM $f$ and humans $g$ on a set of $n$ topics $T=\{t_{1},t_{2},...,t_{n}\}$ by measuring emotions (resp. moral foundations) expressed in tweets about each topic $t_{i}\in T$. Human-written tweets about $t$ are available in a dataset (e.g., COVID-19 Tweets or Roe v. Wade Tweets). To create LM’s tweets about $t_{i}$, we prompt it on the topic to generate a set of $m$ responses $R=\{r_{1},r_{2},...,r_{m}\}$. We compare $\hat{D}(t_{i})$, the distribution of emotions (resp. moral foundations) in LM-generated tweets on topic $t_{i}$, and $D(t_{i})$, the distribution in human-authored tweets on the same topic. We measure affective alignment on a topic $t_{i}$ as $S^{t_{i}}(f,g)\in[0,1]$, using (1 - Jensen-Shannon Distance) between the distributions $\hat{D}(t_{i})$ and $D(t_{i})$. The alignment of LM $f$ with humans $g$ on the set of topics $T$ is averaged over that for each topic $t_{i}$ in it:
@@
Figure 2: Default affect alignment $S$ of different LMs with ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by emotions. For each LM, the alignment is averaged over that on different topics related to the issue, with the means shown by circles and the standard deviations shown by errors bars. Base LMs and instruction-tuned LMs are separated by the black horizontal dashed line. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
@@ 3.1 Representativeness of Affect under Default Prompting
Figures 2 and 5 (in Appendix) report the affective alignment of various LMs with liberals ($g_{l}$) and conservatives ($g_{c}$) in the two datasets. Given that the patterns of alignment measured by emotions and moral sentiments are similar, we focus on the emotional alignment (Figure 2).
Do the models exhibit strong affective alignment? Defining a precise threshold for “strong” alignment is challenging. We consider as baseline the alignment between the two ideological groups, i.e. emotion similarity between liberals and conservatives in online discourses (vertical lines in Figure 2). Any alignment falling short of this benchmark could be deemed insufficient, given the profound divisions in contemporary sociopolitical discourse Rao et al. (2023). This baseline is henceforth referred to as the “partisan alignment baseline”.
From Figure 2, it is evident that all LMs fall short of the partisan alignment baseline, indicating weak alignment. Base LMs, trained on causal language modeling tasks without explicit affective alignment tuning, seem to lack the capacity to learn affect during the pretraining phase. Instruction-tuned models, despite undergoing instruction-based and RLHF training to foster alignment with human values, do not appear to extend this alignment to emotional or moral dimensions. Notably, even sophisticated models like GPT-3.5 exhibit heightened misalignment compared to base models. This could be attributed to the models’ intricate architectures and training processes, which may inadvertently amplify misalignment due to their complexity and sensitivity to the training data’s composition.
@@
Do the models represent each group equitably? Observing Figures 1(a) and 4(a) (in Appendix), it is apparent that on COVID-19 Tweets, all LMs reveal liberal tendencies, as the alignment with liberals is consistently higher. Given the novelty of COVID-19 and its prevalence on social media, where liberal perspectives dominate Shah et al. (2020), we hypothesize that a significant portion of the LMs’ pretraining data is derived from discussions in these forums, thus absorbing emotional and moral tone of liberal narratives.
Conversely, on the Roe v. Wade Tweets (Figure 1(b) and Figure 4(b) in Appendix) the LMs display no discernible political tendencies, with some models exhibiting a slight liberal inclination and others conservative, leading to a generally balanced alignment with both political ideologies. In contrast to COVID-19, Roe v. Wade is a longstanding issue in U.S. history, with discourses extending well beyond social media platforms. Consequently, it is likely that the discussions encompassing both political ideologies are more evenly represented in the pretraining data for LMs.
@@
Figure 3: Steered affective alignment $S$ of different LMs with both ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by emotions, for instruction-tuned LMs. Left-facing triangles represent the models by liberal steered prompting; right-facing triangles represent the models by conservative steered prompting; circles with no filling colors represent the models by default. For each LM, the alignment is averaged over that on different topics detected within the dataset. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.

### v2

@@ Abstract
Language models (LMs) are known to represent the perspectives of some social groups better than others, which may impact their performance, especially on subjective tasks such as content moderation and hate speech detection. To explore how LMs represent different perspectives, existing research focused on positional alignment, i.e., how closely the models mimic the opinions and stances of different groups, e.g., liberals or conservatives. However, human communication also encompasses emotional and moral dimensions. We define the problem of affective alignment, which measures how LMs’ emotional and moral tone represents those of different groups. By comparing the affect of responses generated by 36 LMs to the affect of Twitter messages written by two ideological groups, we observe significant misalignment of LMs with both ideological groups. This misalignment is larger than the partisan divide in the U.S. Even after steering the LMs towards specific ideological perspectives, the misalignment and liberal tendencies of the model persist, suggesting a systemic bias within LMs. [^1] Code and data are available at https://github.com/zihaohe123/llm-affective-alignment.
@@ 1 Introduction
The capacity of language models (LMs) to generate human-like responses to natural language prompts has led to new technologies that support people on cognitive tasks requiring complex judgements. However, researchers found that LMs inherit biases [^2] Throughout this paper, we use “bias” to refer to a systematic statistical tendency, rather than unfairness or prejudice. from humans, as their views are shaped by online users who produced the pretraining data, feedback from crowdworkers during Reinforcement Learning from Human Feedback (RLHF) process Ouyang et al. (2022), and potentially, the decisions made by the model developers themselves Santurkar et al. (2023). In subjective tasks, such as hate speech detection Hartvigsen et al. (2022), content moderation He et al. (2023), and legal judgement Jiang and Yang (2023), these biases may show up as LMs adopting the perspectives of one group while excluding others. This may lead to undesirable downstream consequences, ranging from negative user experiences with LMs to societal-level inequity, division and polarization.
@@
To examine how LMs represent differences in perspectives of different groups, existing research has looked at positional alignment: how closely the model’s opinions or stances mirror those of different social groups Santurkar et al. (2023); Durmus et al. (2023). Using multi-choice survey questions, researchers have demonstrated that language models are misaligned with the US population and represent the perspectives of some demographic groups better than others.
@@
Figure 1: The framework for evaluating affective alignment of LMs. We first prompt LMs to generate tweets on a topic using default prompting or steered prompting. The distributions of emotions and moral sentiments of LM-generated tweets are then compared to that of human-authored tweets. Affective alignment is measured as one minus the Jensen-Shannon distance (JSD) between the two distributions.
However, positional agreement captures just one aspect of alignment. Human communication also carries cues to emotions and moral sentiments–collectively referred to as affect–which are integral to social interaction and cohesion Graham et al. (2009); Iyengar et al. (2019); Mokhberian et al. (2020). While emotions help shape individual’s positions and stances on issues, they are distinct constructs. Take an example of a specific position – a pro-masking stance. The same position can be expressed using contrasting emotions as in the following two statements:
  • Every mask we wear is a badge of honor, showing love and respect for our communities.
  • It’s heartbreaking to see the impact of not wearing masks - lives lost, dreams deferred. Every choice to ditch the mask deepens the crisis. Let’s not add to the pain.
While both statement take a pro-masking stance, the first one expresses positive emotions, and the second negative ones.
@@
How LMs represent affect plays an important role in their performance in downstream tasks, especially subjective tasks. Consider how an LM facilitating online discussions may handle the comment: “Wearing a mask is a personal choice, not a public responsibility.” An LM aligned with conservatives would not flag this comment as it prioritized the moral sentiments of liberty and authority typically associated with conservatives Doğruyol et al. (2019). However, this comment may evoke negative reactions from liberals, as it goes against their deeply-held values of care and fairness, prompting a liberal-aligned LM to flag it. Therefore, we ask:
@@
Whose affect, i.e., emotional and moral tone, do language models reflect?
@@
Our contributions. We define affective alignment as the degree to which a model’s emotional and moral tone matches that of people in similar situations. To study this, we analyze two datasets of Twitter messages on contentious issues like COVID-19 and abortion, disaggregating users by political ideology (liberal or conservative). We prompt 36 LMs of varying sizes to generate statements on topics like “COVID-19 mask mandates” and “abortion rights and access,” then compare the emotions and moral sentiments in the model-generated responses to those of Twitter users from different ideological groups.
@@
We first assess the models’ default affective alignment, based on the responses to default prompts that do not include information about a target demographic (persona). Our findings suggest that LMs show significant misalignment in affect with either ideological group, and the differences are larger than the ideological divide on Twitter. Moreover, consistent with prior findings Santurkar et al. (2023); Perez et al. (2022); Hartmann et al. (2023), all LMs exhibit liberal tendencies on topics related to COVID-19.
@@
Next, we assess LMs’ affective alignment after we provide additional context in the prompt—steer the model—to generate texts from the perspective of liberals or conservatives. The results reveal that steering can help align LMs with the target group’s affect. However, even after steering, the models remain misaligned, and liberal tendencies of LMs cannot be mitigated by steering.
@@
We believe that a deep analysis of the affect expressed by existing LMs is crucial for building AI systems for greater social good. To the best of our knowledge, our work is the first to systematically assess the affective alignment of LMs, which highlights the unequal affective representations of different ideological groups in current LMs. We hope that our framework can help guide future research in better understanding LMs’ representativeness of people from diverse backgrounds on an emotional and moral level. We elaborate the broader impact of the affective alignment in Appendix A.
@@ 2 Measuring Affective Alignment
Figure 1 shows the overall framework. We first prompt LMs to generate tweets on some topic using default prompting or steered prompting. We use state-of-the-art classifiers to measure emotions and moral sentiment of both LM-generated tweets and human-authored tweets. The affective alignment is measured as one minus the Jensen-Shannon distance (JSD) between the distributions of affect.
@@ 2.1 Online Sociopolitical Discourse
To represent human affect empirically, we compile two Twitter datasets on sociopolitical discourse—COVID-19 Tweets and Roe v. Wade Tweets—and identify fine-grained sub-topics related to both issues. We estimate the ideological leanings (liberals vs. conservatives) of users based on the political bias of the news outlets they share. See Appendix B for details.
@@ 2.2 Text Generated by LMs
We categorize LMs into two groups: 1) base LMs Ouyang et al. (2022); Touvron et al. (2023); Almazrouei et al. (2023); Jiang et al. (2023); Conover et al. (2023); Chung et al. (2022); Zheng et al. (2023), that have been pretrained with the causal language modeling objective using data from the Internet; 2) instruction-tuned LMs Ouyang et al. (2022); Touvron et al. (2023); Almazrouei et al. (2023); Jiang et al. (2023); Conover et al. (2023); Chung et al. (2022); Zheng et al. (2023), that have been further finetuned on instructions following and with RLHF for better alignment with human values. To guide the LM to generate open-ended responses on different topics, we use opinion-eliciting prompts. To measure affect representativeness of models in different configurations, we consider the following two settings:
  • Default Prompting – We prompt the LM to generate responses without providing with any additional context. Our goal is to measure the default representativeness of affect expressed by the LM. This would potentially disclose the disparity (if any) of the LM in representing perspectives of different demographic groups. An example prompt for base LMs is “In a tweet about [topic], it’s said that: ”, and for instruction-tuned LMs is “Share your thoughts on [topic] in a tweet.”, where “[topic]” is replaced with the one of the fine-grained topics discovered in the COVID-19 Tweets or the Roe v. Wade Tweets (Table 1 and 2 in Appendix B).
@@
The idea for these two kinds of prompting is inspired by previous works Santurkar et al. (2023); Durmus et al. (2023). To mitigate the effect of the model’s sensitivity to the specific wording in a prompt, we craft 10 different prompts for the base LMs and instruction-tuned LMs, using default prompting and steered prompting, respectively (Table 4 in Appendix). For each fine-grained topic, we generate 2,000 responses, using 2,000 prompts randomly sampled (with replacement) from the 10 candidate prompts. For more details on the generation process, please see Appendix D.1.
@@ 2.3 Measuring Affect
Detecting Emotions. Emotions are a powerful element of human communication vanKleef et al. (2016). To detect emotions, we use SpanEmo Alhuzali et al. (2021), fine-tuned on top of BERT Devlin et al. (2019) on the SemEval 2018 1e-c data Mohammad et al. (2018), which is specifically curated from Twitter and widely recognized as a benchmark for emotion detection on social media. SpanEmo learns the correlations among the emotions and achieves a micro-F1 score of 0.713 on this dataset, outperforming several other baselines and achieving the state-of-the-art in detecting emotions on Twitter data. In addition, Rao et al. (2023a) annotated the emotions of a subset of the Roe v. Wade tweets that we use in our paper, and further evaluated SpanEmo’s performance on it. They showed an average accuracy of over 0.9 across different emotions. We measure the following emotions: anticipation, joy, love, trust, optimism, anger, disgust, fear, sadness, pessimism and surprise. The model returns a score giving the confidence that a tweet expresses an emotion. We average scores over all tweets with that emotion.
Detecting Moral Sentiments. Moral Foundations Theory Haidt et al. (2007) posits that individuals’ moral perspectives are a combination of a set of foundational values. These moral foundations are quantified along five dimensions: dislike of suffering (care/harm), dislike of cheating (fairness/cheating), group loyalty (loyalty/betrayal), respect for authority and tradition (authority/subversion), and concerns with purity and contamination (purity/degradation). These moral dimensions are crucial for understanding the values driving liberal and conservative discourse. We use DAMF Guo et al. (2023b) for moral sentiment detection, which is finetuned on top of BERT Devlin et al. (2019) on three Twitter datasets (including COVID-19 tweets and abortion-related tweets studied in this paper) and one news article dataset. The large amount and the variety of topics in the training data helps mitigate the data distribution shift during inference. Guo et al. (2023b) annotated the moral foundations of a subset of the COVID-19 tweets that we use in our paper, and further evaluated the performance of DAMF on it, which led to F1-score of 0.71 . The model returns a value indicating a confidence that a tweet expresses a moral foundation. We average scores over all tweets with that moral foundation.
@@ 2.4 Measuring Alignment
Let us represent an LM as $f$ and a group of humans as $g$. We aim to measure affective alignment $S^{T}(f,g)$ between the LM $f$ and humans $g$ on a set of $n$ topics $T=\{t_{1},t_{2},...,t_{n}\}$ by measuring emotions (resp. moral sentiments) expressed in tweets about each topic $t_{i}\in T$. Human-authored tweets about $t$ are available in a dataset (e.g., COVID-19 Tweets or Roe v. Wade Tweets). To create LM’s tweets about $t_{i}$, we prompt it on the topic to generate a set of $m$ responses $R=\{r_{1},r_{2},...,r_{m}\}$. We compare $\hat{D}(t_{i})$, the distribution of emotions (resp. moral foundations) in LM-generated tweets on topic $t_{i}$, and $D(t_{i})$, the distribution in human-authored tweets on the same topic. We measure affective alignment on a topic $t_{i}$ as $S^{t_{i}}(f,g)\in[0,1]$, using (1 - Jensen-Shannon Distance) between the distributions $\hat{D}(t_{i})$ and $D(t_{i})$. The alignment of LM $f$ with humans $g$ on the set of topics $T$ is averaged over that for each topic $t_{i}$ in it:
@@
2.4.1 Proximity Between Emotions
To more accurately capture the interrelated nature of emotions [^3] We only consider the agreement between different emotions but not moral foundations, as there is no existing work on modeling the structure of moral foundations as the Plutchik’s wheel for emotions., we integrate the concept of emotional proximity using the Plutchik Emotion Agreement (PEA). For instance, emotions like joy and love are closer on the emotional spectrum than joy and anger. To quantitatively capture these relationships, we utilize the Plutchik Emotion Agreement (PEA) Desai et al. (2020), leveraging the Plutchik wheel Plutchik (2001) which organizes emotions into a spatial model indicative of their relational proximity. The PEA is calculated from the polar coordinates of each emotion on the wheel, as
$\text{PEA}(e_{i},e_{j})=max(|1-\frac{1}{\pi}|f(e_{i}),f(e_{j})||),$ (2)
where $e_{i}$ and $e_{j}$ are two different emotions, and $f(e)$ represents the polar coordinate of the emotion. The emotion proximity matrix $A$, where $A_{ij}$ represents the proximity between $e_{i}$ and $e_{j}$, is shown in Appendix C. This methodological adjustment allows us to account for the interconnected nature of emotional expressions, refining our alignment measurements. We weight the emotion distributions vectors using the proximity matrix, as $\hat{D}^{\prime}(t)=A\cdot\hat{D}(t)$, and $D^{\prime}(t)=A\cdot D(t)$, then use the weighted vectors for computing the affective alignment as in Equation 1.
@@ 2.4.1 Proximity Between Emotions
Figure 2: Default affect alignment $S$ of different LMs with ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by emotions. * indicates that the alignment of the model with both ideological groups are significantly different at $p<0.05$. For each LM, the alignment is averaged over that on different topics related to the issue, with the means shown by circles and the standard deviations shown by errors bars. Base LMs and instruction-tuned LMs are separated by the black horizontal dashed line. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
@@ 3.1 Representativeness of Affect under Default Prompting
Figures 2 reports the affective alignment of various LMs with liberals ($g_{l}$) and conservatives ($g_{c}$) by default prompting in the two datasets, measured by emotions. Please refer to Appendix D.2 for results measured by moral sentiments. Since the patterns of alignment measured by emotions and moral sentiments are similar, we focus on the emotional alignment.
Do the models exhibit strong affective alignment? Defining a precise threshold for “strong” alignment is challenging. We consider as baseline the alignment between the two ideological groups on Twitter, i.e. emotion similarity between liberals and conservatives in online discourses (vertical lines in Figure 2). Any alignment falling short of this benchmark could be deemed insufficient, given the profound divisions in contemporary sociopolitical discourse Rao et al. (2023b). This baseline is henceforth referred to as the “partisan alignment baseline”.
From Figure 2, it is evident that nearly all LMs fall short of the partisan alignment baseline, indicating weak alignment. Base LMs, trained on causal language modeling tasks without explicit affective alignment tuning, seem to lack the capacity to learn affect during the pretraining phase. Instruction-tuned models, despite undergoing instruction-based and RLHF training to foster alignment with human values, do not appear to extend this alignment to emotional or moral dimensions. Notably, even sophisticated models like GPT-3.5 exhibit heightened misalignment compared to base models. This could be attributed to the models’ intricate architectures and training processes, which may inadvertently amplify misalignment due to their complexity and sensitivity to the training data’s composition.
@@
Do the models represent each group equitably? It is apparent that on COVID-19, all LMs reveal liberal tendencies, as the alignment with liberals is consistently higher, and the partisan alignment difference is statistically significant. Given the novelty of COVID-19 and its prevalence on social media, where liberal perspectives dominate Shah et al. (2020), we hypothesize that a significant portion of the LMs’ pretraining data is derived from discussions in these forums, and thus LMs absorb more emotional and moral tone of liberal narratives.
Conversely, on the Roe v. Wade Tweets (Figure 2(b)) the LMs display no discernible political tendencies, with some models exhibiting a slight liberal inclination and others conservative, leading to a generally balanced alignment with both political ideologies. In fact, the partisan alignment difference is not statistically significant on Roe v. Wade. In contrast to COVID-19, Roe v. Wade is a longstanding issue in U.S. history, with discourses extending well beyond social media platforms. Consequently, it is likely that the discussions encompassing both political ideologies are more evenly represented in the pretraining data for LMs.
@@
Figure 3: Steered affective alignment $S$ of different instruction-tuned LMs with both ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by emotions, for . * indicates that the alignment of the liberal steered model with both ideological groups are significantly different at $p<0.05$; ^ indicates that for the conservative steered model. Left-facing triangles represent the models by liberal steered prompting; right-facing triangles represent the models by conservative steered prompting; circles with no filling colors represent the models by default. For each LM, the alignment is averaged over that on different topics detected within the dataset. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.